# U-DECN: End-to-End Underwater Object Detection ConvNet with Improved DeNoising Training

### v1

@@ Abstract
Underwater object detection has higher requirements of running speed and deployment efficiency for the detector due to its specific environmental challenges. NMS of two- or one-stage object detectors and transformer architecture of query-based end-to-end object detectors are not conducive to deployment on underwater embedded devices with limited processing power. As for the detrimental effect of underwater color cast noise, recent underwater object detectors make network architecture or training complex, which also hinders their application and deployment on underwater vehicle platforms. In this paper, we propose the Underwater DECO with improved deNoising training (U-DECN), the query-based end-to-end object detector (with ConvNet encoder-decoder architecture) for underwater color cast noise that addresses the above problems. We integrate advanced technologies from DETR variants into DECO and design optimization methods specifically for the ConvNet architecture, including Separate Contrastive DeNoising Forward and Deformable Convolution in SIM. To address the underwater color cast noise issue, we propose an underwater color denoising query to improve the generalization of the model for the biased object feature information by different color cast noise. Our U-DECN, with ResNet-50 backbone, achieves 61.4 AP (50 epochs), 63.3 AP (72 epochs), 64.0 AP (100 epochs) on DUO, and 21 FPS (5 times faster than Deformable DETR and DINO 4 FPS) on NVIDIA AGX Orin by TensorRT FP16, outperforming the other state-of-the-art query-based end-to-end object detectors. The code is available at https://github.com/LEFTeyex/U-DECN.
@@ 1 Introduction
Object detection is a fundamental task in computer vision with widespread real-world applications. Among numerous applications, underwater object detection has become a unique and important research field due to its specific environmental challenges. In underwater environments, the absorption and scattering of light by water media lead to problems such as color cast and low contrast in underwater images, which seriously reduce the performance of underwater object detection, especially the significant effect of color cast [26]. In addition, underwater object detectors are typically deployed on embedded devices with limited processing power and required real-time processing. Therefore, underwater object detection not only needs to overcome the effect of underwater color cast noise but also needs to have high running speed and deployment efficiency.
Figure 1: Underwater DECN architecture for color cast noise.
Object detection has been developed rapidly in the era of deep learning. Typical two-stage object detectors like Faster R-CNN [23] and SPPNet [8], etc. have achieved high performance by refining priori bounding box predictions. Meanwhile, one-stage detectors like SSD [16], RetinaNet [13], and YOLO series [22, 21], etc. simplify the detection pipeline by directly predicting the objects, improving the speed of object detection. Most of the above object detectors are built upon convolutional neural networks (CNNs or ConvNets) and typically the non-maximum suppression (NMS) strategy is utilized for post-processing to remove duplicate detection results. The computational process of the NMS strategy involves sorting and comparison operations that increase computation time and resource consumption, reducing model deployment efficiency, especially on embedded devices with limited processing power.
While the Detection Transformer (DETR [1]) and its variants [31, 28] have shown prominent performance on object detection tasks. DETR refactors the object detection pipeline as a set prediction problem and directly obtains a fixed set of objects via a transformer encoder-decoder architecture. This design enables object detectors like DETR to eliminate the complicated NMS post-processing, resulting in a query-based end-to-end object detection pipeline. However, DETR and its variants contain operations of attention modules that AI chips typically cannot support well. It increases the difficulty of deploying such object detectors and reduces their running speed and deployment efficiency.
@@
In the field of underwater object detection, some works [18, 10] address the problem of underwater color cast noise by using deep learning image enhancement methods for underwater image denoising preprocessing. The image enhancement network module and the detector are jointly trained to improve the interaction between them. Other works improve the generalization of object detection models against underwater color cast noise by domain generalization [5] or multi-noise model fusion [3]. However, the above underwater object detection work makes the network architecture or training of the detector complex. Most of them contain NMS post-processing, which hinders the application and deployment of underwater object detectors.
In this paper, we design a new DECO-like underwater object detector based on DECO [4] and DINO [28]. It is named as U-DECN (Underwater DECO with improved deNoising training), its simplified architecture is shown in Fig. 1. To improve both convergence speed and performance, we integrate mixed query selection, look forward twice, contrastive denoising training, hybrid encoder methods to DECO. Meanwhile, we propose the separate contrastive denoising forward method to make the ConvNet encoder-decoder architecture capable of performing contrastive denoising training. We integrate the deformable convolution into the query self-interaction module in the decoder, in order to correct the position of initialized (from the encoder) positional queries due to the mixed query selection method. To address the underwater color cast noise issue, we propose an underwater color denoising query to improve the generalization of the model for the biased object feature information by different color cast noise.
U-DECN achieves an ideal trade-off between running speed (deployment efficiency) and accuracy. Specifically, with ResNet-50 [9] backbone, U-DECN achieves 61.4 AP (50 epochs), 63.3 AP (72 epochs), 64.0 AP (100 epochs) on DUO [14], 29 FPS on an NVIDIA A6000 GPU, and 21 FPS (5 times faster) on NVIDIA AGX Orin by TensorRT FP16, outperforming the other state-of-the-art query-based end-to-end object detectors. Extensive ablation studies are also conducted to validate the effectiveness of our methods in improving both convergence speed and performance.
@@
  • We design a new end-to-end DECO-like underwater object detector called U-DECN, with several techniques, Two-Stage Bounding Box Refinement, Separate Contrastive DeNoising Forward, Deformable Convolution in SIM, and Hybrid Encoder for different parts of U-DECN.
@@
  • We propose an underwater color denoising query to improve the generalization for underwater color cast noise.
  • Our U-DECN is extensively evaluated on the DUO benchmark which obtains competitive in terms of running speed (deployment efficiency) and accuracy.
The rest of this paper is organized as follows: Sec. 2 reviews some related works. Sec. 3 describes the details of the proposed U-DECN. Sec. 4 contains relevant experiments and analysis. Sec. 5 is about conclusion.
@@ 2 Related Work
Two- and One-Stage Object Detectors. Classical object detectors based on ConvNets are either two-stage or one-stage models. Two-stage models, such as Faster R-CNN [23]and SPPNet [8], usually use a region proposal network (RPN) to propose priori bounding boxes, which are then refined in the second stage. Although they perform well, the two-stage bounding box refinement reduces their running speed. One-stage models such as SSD [16], RetinaNet [13], and YOLO series [22, 21] directly output offsets relative to predefined anchors, improving the speed of object detection. Moreover, most of the above object detectors need the NMS post-processing to remove duplicate detection results, which reduces model deployment efficiency.
@@
Query-Based End-to-End Object Detectors. The pioneering work DETR [1] utilizes a transformer encoder-decoder architecture and models object detection as a set prediction problem. It directly predicts a fixed number of objects and eliminates the hand-crafted anchor and NMS components. Many follow-up studies have attempted to address the slow training convergence, lack of multi-scale features, and lower performance issues of DETR. For instance, Deformable DETR [31] only attends to certain sampling points around a reference point by introducing predicted 2D anchor points and multi-scale deformable attention. DAB-DETR [15] further extends 2D anchor points to 4D anchor box coordinates to represent queries and dynamically update boxes in each decoder layer. DN-DETR [11] and DINO [28] introduce query denoising training and mixed query selection methods to improve DETR training speed and performance. RT-DETR [30] design efficient hybrid encoder architecture and IoU-aware query selection method to improve the efficiency of multi-scale feature interaction and the quality of initialized queries. Recently, a query-based end-to-end object detection framework with ConvNets named DECO [4] has been proposed. DECO is composed of a convolutional encoder-decoder architecture, rather than a sophisticated transformer architecture, which performs the interaction between object queries and image features via convolutional self-interaction module (SIM) and cross-interaction module (CIM) in decoder layers. It improves the running speed and deployment efficiency of the query-based end-to-end object detector, but it still has some issues to solve, like DETR.
Underwater Object Detectors. In underwater environments, the effect of underwater environmental noise on object detection performance is particularly severe, especially color cast noise [26]. Some traditional image enhancement methods [29, 7] are used in object detection to pre-process the input image. The follow-up ULO [24] designs a module for the dynamic prediction of hyper-parameters in the image enhancement method for each image. Several similar works use image enhancement methods based on deep learning with joint training to improve detection performance under color cast noise, such as GAN [24], CNN [10], etc. Another kind of method mainly focuses on the generalization of the object detection model in the condition that the input image contains noise. DMCL [5] optimizes the training method and applies contrastive learning, which is designed to improve the domain generalization of the detector. FERNet [6] proposes a receptive field enhancement module for the backbone to exploit multi-scale underwater semantic features. SWIPENET [3] designs a sample-weighted hyper network and a robust training paradigm to learn potential information from different noises. The complexity of network architecture and training in underwater object detection, along with NMS post-processing, hinders the application and deployment of underwater object detectors.
3 U-DECN
3.1 Preliminaries
DETR [1] and its variants [31, 28] based on transformer architecture have attracted great research interest due to the good accuracy and runtime performance for object detection. Although transformers have shown great power in computer vision tasks like image classification, object detection, segmentation etc., some recent works reveal the potential of ConvNet-based architecture as the common backbone, e.g., ConvNeXt [17] and ConvFormer [27].
Detection ConvNet (DECO [4]) designs the ConvNet encoder-decoder architecture for object detection and utilizes the same prediction loss as in DETR, which uses bipartite matching to find paired predicted and ground truth objects. The DECO encoder is built upon ConvNeXt blocks. Due to the permutation-variant property of ConvNet architecture, its encoder layers could get rid of any positional embedding. Similarly to DETR, the DECO decoder is stacked with two modules in each layer: a self-interaction module (SIM) and a cross-interaction module (CIM) based on ConvNets. The SIM is stacked with a $9\times 9$ large kernel depth-wise convolution and $1\times 1$ convolutions to perform long-range perceptual feature extraction for object queries. Given $N$ object queries $\bm{q}\in\mathbb{R}^{N\times d}$ (belonging to the content query, where $d$ is the embedding dimension), DECO reshapes the queries to $\mathbb{R}^{d\times h_{q}\times w_{q}}$ and feeds them into convolutional layers. While $N=h_{q}\times w_{q}$ indicates that the setting of the $N$ value is limited. In addition, DECO appends positional queries in the SIM part, which are processed in the same way as content queries, and both its content and positional queries are learnable. The image feature embeddings ($\bm{x}_{e}\in\mathbb{R}^{d\times h\times w}$) from the encoder output and the object query embeddings ($\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$) produced from the SIM are interacted with by the CIM. The object queries $\bm{q}$ are upsampled first to obtain $\hat{\bm{q}}\in\mathbb{R}^{d\times h\times w}$ so that it has the same size as the image feature $\bm{x}_{e}$. Then the upsampled object queries and the image feature embeddings are fused using add operations, followed by a $9\times 9$ large kernel depth-wise convolution to allow the object queries to capture the spatial information from the image feature. After some residual connection and feed-forward network (FFN) operations, an adaptive max-pooling layer is utilized to downsample the object queries back to the size of $\mathbb{R}^{d\times h_{q}\times w_{q}}$, which are further processed by the following decoder layers. In this work, we will further explore the compatibility of DECO with more advanced techniques in DETR variants and improve its generalization against underwater color cast noise, the framework of our proposed U-DECN is shown in Fig. 2.
@@ 3.1 Preliminaries
3.2 Two-Stage Bounding Box Refinement
@@
The mix of dynamically selected positional queries and static content queries achieves great success in the transformer encoder-decoder architecture [28, 30]. The attention module of the transformer has a permutation-invariant property that utilizes selected image features to construct positional queries that can focus on the feature information of their location. The ConvNet encoder-decoder architecture must first reshape the object queries to $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$, then upsample the queries to the same size as the image feature by bilinear interpolation, and finally perform add operations and convolutional interactions with the image feature. However, convolutional interaction requires that the location of the object queries be roughly aligned with the image features being queried to achieve better interaction. The positional queries, which are randomly initialized in DECO, can achieve positional alignment with image features by learning. But the dynamically selected positional queries are uncertainly arranged in $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$, it is difficult to align them with the image features by the $9\times 9$ large kernel depth-wise convolution in SIM. We named it the query feature misalignment issue. This issue makes mixed query selection limited in the ConvNet encoder-decoder architecture, details in Sec. 3.4. Meanwhile, the ConvNet encoder and decoder can only process single-level features and cannot simultaneously process multi-level features like the detector transformer [31], resulting in a lack of multi-scale feature information, details in Sec. 3.5.

### v2

@@ Abstract
Underwater object detection has higher requirements of running speed and deployment efficiency for the detector due to its specific environmental challenges. NMS of two- or one-stage object detectors and transformer architecture of query-based end-to-end object detectors are not conducive to deployment on underwater embedded devices with limited processing power. As for the detrimental effect of underwater color cast noise, recent underwater object detectors make network architecture or training complex, which also hinders their application and deployment on unmanned underwater vehicles. In this paper, we propose the Underwater DECO with improved deNoising training (U-DECN), the query-based end-to-end object detector (with ConvNet encoder-decoder architecture) for underwater color cast noise that addresses the above problems. We integrate advanced technologies from DETR variants into DECO and design optimization methods specifically for the ConvNet architecture, including Deformable Convolution in SIM and Separate Contrastive DeNoising Forward methods. To address the underwater color cast noise issue, we propose an Underwater Color DeNoising Query method to improve the generalization of the model for the biased object feature information by different color cast noise. Our U-DECN, with ResNet-50 backbone, achieves the best 64.0 AP on DUO and the best 58.1 AP on RUOD, and 21 FPS (5 times faster than Deformable DETR and DINO 4 FPS) on NVIDIA AGX Orin by TensorRT FP16, outperforming the other state-of-the-art query-based end-to-end object detectors. The code is available at https://github.com/LEFTeyex/U-DECN.
@@ 1 Introduction
Approximately 71% of the Earth’s surface is covered by oceans, presenting vast opportunities for exploration. However, the hyperbaric and hypoxic conditions of the deep-sea environment pose significant challenges to direct human exploration. Recent advancements in unmanned underwater vehicles (UUVs) and deep learning, particularly in underwater optical image object detection, have significantly enhanced the capabilities of ocean exploration. In underwater environments, the absorption and scattering of light by water media lead to problems such as color cast in underwater images, which seriously reduce the performance of underwater object detection [29, 23, 32]. In addition, underwater object detectors are typically deployed on embedded devices with limited processing power and required real-time processing. Therefore, underwater object detection not only needs to overcome the effect of underwater color cast noise but also needs to have high running speed and deployment efficiency.
Figure 1: U-DECN architecture for color cast noise.
Object detection has been developed rapidly in the era of deep learning. Typical two-stage object detectors like Faster R-CNN [26] and SPPNet [10], etc. have achieved high performance by refining priori bounding box predictions. Meanwhile, one-stage detectors like SSD [19], RetinaNet [16], and YOLO series [25, 24], etc. simplify the detection pipeline by directly predicting the objects, improving the speed of object detection. Most of the above object detectors are built upon convolutional neural networks (CNNs or ConvNets) and typically the non-maximum suppression (NMS) strategy is utilized for post-processing to remove duplicate detection results. The computational process of the NMS strategy involves sorting and comparison operations that increase computation time and resource consumption, reducing model deployment efficiency, especially on embedded devices with limited processing power. The Detection Transformer (DETR [1]) and its variants [34, 30] have shown prominent performance on object detection tasks. DETR refactors the object detection pipeline as a set prediction problem and directly obtains a fixed set of objects via a transformer encoder-decoder architecture. This design enables object detectors like DETR to eliminate the complicated NMS post-processing, resulting in a query-based end-to-end object detection pipeline. However, DETR and its variants contain operations of attention modules that AI chips typically cannot support well. It increases the difficulty of deploying such object detectors and reduces their running speed and deployment efficiency.
@@
In the field of underwater object detection, the presence of underwater environmental noise, especially color cast, severely degrades the quality of captured images and poses significant challenges to robust detection [29]. To address this, some works [31, 9, 14] use traditional image enhancement methods as a preprocessing step. More recent works adopt deep learning image enhancement techniques [20, 12], where the enhancement network and the detector are jointly trained to improve feature interaction. For instance, ULO [27] introduces a hyper-parameter prediction module that dynamically adjusts enhancement strategies per image, and similar approaches using GAN or CNN aim to suppress the impact of color cast noise during detection. Another kind of research focuses on improving the generalization of object detectors under noisy conditions. Domain generalization techniques [2], contrastive learning strategies such as DMCL [6], and multi-noise model fusion approaches like SWIPENET [4] attempt to increase model robustness to varied underwater conditions. FERNet [7] enhances multi-scale semantic representation via receptive field expansion, while SWIPENET designs a sample-weighted hyper network and robust training paradigms to handle noisy inputs. However, many of these methods rely on complex architectures or training procedures and still require non-maximum suppression (NMS) as a post-processing step. The computational complexity introduced by NMS and intricate model components significantly hinders the application and deployment of underwater object detectors.
In this paper, we design a new DECO-like underwater object detector based on DECO [5] and DINO [30]. It is named as U-DECN (Underwater DECO with improved deNoising training), its simplified architecture is shown in Fig. 1. To introduce multi-scale feature information, we append Hybrid Encoder to DECO. Meanwhile, introducing Two-Stage Bounding Box Refinement to improve model performance and to address query feature misalignment issue caused by Two-Stage Bounding Box Refinement, we design Deformable Convolution in SIM method to correct the position of initialized (from the encoder) positional queries. To address the underwater color cast noise issue, we propose an Underwater Color DeNoising Query method to improve the generalization of the model for the biased object feature information by different color cast noise. Finally, we propose the Separate Contrastive DeNoising Forward method to make the ConvNet encoder-decoder architecture capable of performing contrastive denoising training, which is to improve the convergence speed of training.
U-DECN achieves an ideal trade-off between running speed (deployment efficiency) and accuracy. Specifically, with the ResNet-50 [11] backbone, U-DECN achieves the best 64.0 AP on DUO [17] and the best 58.1 AP on RUOD [8], 29 FPS on an NVIDIA A6000 GPU, and 21 FPS (5 times faster) on NVIDIA AGX Orin by TensorRT FP16, outperforming the other state-of-the-art query-based end-to-end object detectors. Extensive ablation studies are also conducted to validate the effectiveness of our methods in improving both convergence speed and performance.
@@
  • We design a new end-to-end DECO-like underwater object detector called U-DECN, with several techniques, Hybrid Encoder (introducing multi-scale feature information), Two-Stage Bounding Box Refinement (improving model performance), Deformable Convolution in SIM (addressing query feature misalignment issue caused by Two-Stage Bounding Box Refinement), Underwater Color DeNoising Query (improving the generalization for underwater color cast noise), and Separate Contrastive DeNoising Forward (improving the convergence speed of training) for different parts of U-DECN.
  • Our U-DECN is extensively evaluated on the DUO and RUOD benchmark which obtains competitive in terms of running speed, deployment efficiency, and accuracy. Ablation studies demonstrate the effectiveness of our proposed methods.
The rest of this paper is organized as follows: Section 2 reviews some related works. Section 3 describes the details of the proposed U-DECN. Section 4 contains relevant experiments and analysis. Section 5 is about conclusion.
@@ 2 Related Work
2.1 Transformer-Based End-to-End Object Detectors
@@
The pioneering work DETR [1] utilizes a transformer encoder-decoder architecture and models object detection as a set prediction problem. It directly predicts a fixed number of objects and eliminates the hand-crafted anchor and NMS components. Many follow-up studies have attempted to address the slow training convergence, lack of multi-scale features, and lower performance issues of DETR. For instance, Deformable DETR [34] only attends to certain sampling points around a reference point by introducing predicted 2D anchor points and multi-scale deformable attention. DAB-DETR [18] further extends 2D anchor points to 4D anchor box coordinates to represent queries and dynamically update boxes in each decoder layer. DN-DETR [13] and DINO [30] introduce query denoising training and mixed query selection methods to improve DETR training speed and performance. RT-DETR [33] design efficient hybrid encoder architecture and IoU-aware query selection method to improve the efficiency of multi-scale feature interaction and the quality of initialized queries.
@@
2.2 ConvNet-Based End-to-End Object Detectors
Recently, a ConvNet-based end-to-end object detection framework named DECO [5] has been proposed. DECO is composed of a convolutional encoder-decoder architecture, rather than a sophisticated transformer architecture, which performs the interaction between object queries and image features via convolutional self-interaction module (SIM) and cross-interaction module (CIM) in decoder layers. Detection ConvNet (DECO) designs the ConvNet encoder-decoder architecture for object detection and utilizes the same prediction loss as in DETR, which uses bipartite matching to find paired predicted and ground truth objects. The DECO encoder is built upon ConvNeXt blocks. Due to the permutation-variant property of ConvNet architecture, its encoder layers could get rid of any positional embedding. Similarly to DETR, the DECO decoder is stacked with two modules in each layer: a self-interaction module (SIM) and a cross-interaction module (CIM) based on ConvNets. The SIM is stacked with a $9\times 9$ large kernel depth-wise convolution and $1\times 1$ convolutions to perform long-range perceptual feature extraction for object queries. Given $N$ object queries $\bm{q}\in\mathbb{R}^{N\times d}$ (belonging to the content query, where $d$ is the embedding dimension), DECO reshapes the queries to $\mathbb{R}^{d\times h_{q}\times w_{q}}$ and feeds them into convolutional layers. While $N=h_{q}\times w_{q}$ indicates that the setting of the $N$ value is limited. In addition, DECO appends positional queries in the SIM part, which are processed in the same way as content queries, and both its content and positional queries are learnable. The image feature embeddings ($\bm{x}_{e}\in\mathbb{R}^{d\times h\times w}$) from the encoder output and the object query embeddings ($\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$) produced from the SIM are interacted with by the CIM. The object queries $\bm{q}$ are upsampled first to obtain $\hat{\bm{q}}\in\mathbb{R}^{d\times h\times w}$ so that it has the same size as the image feature $\bm{x}_{e}$. Then the upsampled object queries and the image feature embeddings are fused using add operations, followed by a $9\times 9$ large kernel depth-wise convolution to allow the object queries to capture the spatial information from the image feature. After some residual connection and feed-forward network (FFN) operations, an adaptive max-pooling layer is utilized to downsample the object queries back to the size of $\mathbb{R}^{d\times h_{q}\times w_{q}}$, which are further processed by the following decoder layers. It improves the running speed and deployment efficiency of the end-to-end object detector, but it still has some issues to solve, like DETR.
3 Method
In this work, we will further explore the compatibility of DECO [5] with more advanced techniques in DETR variants [34, 30] and improve its generalization against underwater color cast noise, the framework of our proposed U-DECN is shown in Fig. 2.
3.1 Hybrid Encoder
The ConvNet encoder-decoder architecture of DECO [5] can only process single-level features, and cannot perform the interaction of multi-level features like the detector transformer [34], resulting in a lack of multi-scale feature information and reducing the performance of the detector for objects on different scales. The simultaneous intra-scale and cross-scale feature interaction in the transformer encoder has been proven to be inefficient cf. RT-DETR [33], resulting in a computational bottleneck. RT-DETR proposes an efficient hybrid encoder, consisting of the Attention-based Intra-scale Feature Interaction (AIFI) and the CNN-based Cross-scale Feature Fusion (CCFF) modules. Specifically, AIFI further reduces the computational cost by performing the intra-scale interaction only on the highest-level feature. Then, CCFF fuses the highest-level feature and other low-level features and outputs the fusion features of the corresponding levels. To introduce multi-scale feature information, we also use the hybrid encoder architecture, as shown in Fig. 2, where AIFI i.e. our ConvNet encoder. We take the highest-level feature of CCFF for decoder, since our ConvNet decoder could still only process single-level features. Detailed CCFF structure reference RT-DETR.
3.2 Two-Stage Bounding Box Refinement and Deformable Convolution in SIM
@@
The mix of dynamically selected positional queries and static content queries achieves great success in the transformer encoder-decoder architecture [30, 33]. The attention module of the transformer has a permutation-invariant property that utilizes selected image features to construct positional queries that can focus on the feature information of their location. The ConvNet encoder-decoder architecture must first reshape the object queries to $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$, then upsample the queries to the same size as the image feature by bilinear interpolation, and finally perform add operations and convolutional interactions with the image feature. However, convolutional interaction requires that the location of the object queries be roughly aligned with the image features being queried to achieve better interaction. The positional queries, which are randomly initialized in DECO, can achieve positional alignment with image features by learning. But the dynamically selected positional queries are uncertainly arranged in $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$, it is difficult to align them with the image features by the $9\times 9$ large kernel depth-wise convolution in SIM. We named it the query feature misalignment issue. This issue makes mixed query selection limited in the ConvNet encoder-decoder architecture.
As mentioned above, the positional queries constructed using selected top-K features have the query feature misalignment issue. In CIM, due to the strong positional inductive bias of ConvNets, queries $\bm{q}\in\mathbb{R}^{d\times h_{q}\times w_{q}}$ could only interact with the encoder feature corresponding to their tensor positions. The misalignment between uncertainly arranged positional queries and encoder features results in low interaction efficiency between them, which in turn reduces training convergence speed and model performance in Section 4.4.1. In SIM, even if it uses a $9\times 9$ large kernel depth-wise convolution for self-interaction between queries, which has a large receptive field, it is difficult to correct the position of uncertainly arranged queries. To address this, we replace the $9\times 9$ large kernel depth-wise convolution in SIM with a $9\times 9$ large kernel deformable convolution (DCNv3 [28]), as shown in Fig. 3. The deformable convolution achieves deformable operations by predicting the offset of each feature point in the feature map, and the predicted feature point offset is equivalent to using self-attention to correct the feature position. Based on the above analysis, deformable convolution can also correct the position of uncertainly arranged positional queries in tensor $\bm{q}$ by learning the bounding box position information in priori positional queries to improve the efficiency of interaction with encoder features.
3.3 Underwater Color DeNoising Query
We rethink the effect of underwater color cast noise and consider the color cast noise as color fog. However, there is a difference between this color fog and normal white fog. White fog on land blocks the light reflected from objects, resulting in a lack of object feature information. Color fog is caused by the absorption of light by water media. Although the light reflected from objects is also absorbed causing color cast, the color cast results in a bias rather than a lack of object feature information. Therefore, in terms of noise generalization, we propose an underwater Color DeNoising (ColorDN) query to improve the generalization of the model for the biased object feature information by different color cast noise priori, as shown in Figs. 1 and 2. Underwater color cast noise can cause the overall hue of underwater images to appear blue, green, etc., so the mean of RGB channel of the image can roughly reflect the color cast of underwater images. We define the mean of the three RGB channels in the image as the priori RGB values of color fog in underwater images, and encode it as a ColorDN query using a linear layer, which is as follows:
$\bm{q}_{colordn}=\mathrm{Linear}\left(\mu_{c}(\bm{x}_{input})\right)$ (1)
$\bm{q}_{content}=\bm{q}_{content}+\bm{q}_{colordn}$ (2)
where for different content queries (including the CDN content queries in Section 3.4), we use the tensor repeat operation to extend the ColorDN query to make their quantity consistent. We introduce priori color cast noise information into content queries as the prompt of underwater color cast noise to enable the model to query the corresponding noise information, thereby generalizing the bias of object feature information. The underwater ColorDN query improves the generalization of the detector to color cast noise in a simple and implicit way.